OpenStudy (k8lyn911):

Calculus II Area between curves homework Sketch the region enclosed by the given curves. Find its area. y = 4 cos 4x, y = 4 - 4 cos 4x, from 0 to pi/4 I've sketched the graph, but I'm not sure how to calculate the exact point of intersection. The intersection is (0.26179939, 2), but I don't know what that would be as a fraction.

OpenStudy (anonymous):

you dont have to i don't think

OpenStudy (anonymous):

from 0 to pi/4

OpenStudy (anonymous):

they are telling you the limits of integration so it is not a region enclosed by the two curve, which is probably what you did in another problem just between those endpoints

OpenStudy (anonymous):

oops i see the problem they switch places damn

OpenStudy (k8lyn911):

Yep. There's a point where they cross, and I have to switch the upper and lower functions at that point. The website we do our homework on doesn't like decimals, though, so I have to use fractions to get exact answers.

OpenStudy (anonymous):

think they intersect at \(\frac{\pi}{12}\)which is a drag`

OpenStudy (anonymous):

it is easier to solve without the machine \[4\cos(4x)=4-4\cos(4x)\\ 8\cos(4x)=4\\ \cos(4x)=\frac{1}{2}\]

OpenStudy (anonymous):

unfortunately that makes \[\cos(4x)=\frac{\pi}{3}\]so \[x=\frac{\pi}{12}\]

OpenStudy (k8lyn911):

Oh, so that's how you figure it out. I've got it now. :) Thank you so much!

OpenStudy (anonymous):

yw